### v1

@@ 8.3 Tight Results In The Abelian case
This simple randomized algorithm recovers state of the art results for AMM achieved using linear sketching techniques. Although the error is the same, there might be other qualitative differences. An important difference in the context of deep learning, is the rank of the output matrix. Our empiric trials suggest that the rank of the output matrices in this approach is generally full, even for small $r$. Compared with sketching techniques, where $r$ is also the upper bound on the rank, this is a major difference.
9 Fourier Domain Truncation
The third approach we will present for solving 6.1 is based on the following simple idea: Given $r$, choose (randomly or deterministically) a set $K\subset\hat{G}$ (where $G$ is an abelian group) of $r$ frequencies of the Fourier transform of the convolution c=abc=a*b, compute $c^{\prime}:=\mathcal{F}^{-1}_{G}(\hat{c}|_{K})$, and decode $c^{\prime}$, i.e., return $I_{C}^{\top}c^{\prime}$ where $I_{C}$ is the indexing matrix for $\mathbf{c}$.
In order to show this is a useful approach we must show two things:
  1. 1.
    Bounded Error: We need to bound $\left\lVert I_{C}(c^{\prime})-I_{C}(c)\right\rVert_{2}^{2}$ in terms of $\left\lVert a\right\rVert_{2},\left\lVert b\right\rVert_{2}$ and $r$.
  2. 2.
    Low Complexity: We need to show that $\hat{a}|_{K},\hat{b}|_{K}$ and $c^{\prime}$ can be computed in time $\tilde{O}(r)$.
Remark 9.1.
A trivial bound on the runtime complexity of computing $\hat{a}|_{K},\hat{b}|_{K}$ and $c^{\prime}$ is $O(rkn^{2})$. This follows from observation that the algorithm can be written down as follows: Given inputs $\mathbf{a},\mathbf{b}\in\mathbb{R}^{kn^{2}}$ and a set of $r$ frequencies $K\subset\hat{G}$, computing $\mathbf{c}^{\prime}$ is the same as computing
$f(\mathbf{a},\mathbf{b})=I_{C}^{\top}\left(\mathcal{F}_{G}^{-1}\left(P_{K}([\mathcal{F}_{G}(I_{A}\mathbf{a})]\odot[\mathcal{F}_{G}(I_{B}\mathbf{b})]\right)\right),$
where $P_{K}$ is the idempotent projection onto $K$ (meaning $(P_{K}x)_{\chi}=x_{\chi}\cdot\mathbf{1}_{\chi\in K}$). Note that $P_{K}(x\odot y)=P_{K}(x)\odot P_{K}(y)$, and so by pre-computing some of the matrix products (which are fixed), we obtain smaller matrices and vectors, leading the complexity to $O(rkn^{2})$.
In order to achieve low complexity, we will thus need $K$ to admit some extra structure. We will see examples for this later in the section.
9.1 Basic Properties
Let c=abc=a*b denote the convolved signal and $c^{\prime}=\mathcal{F}^{-1}(\hat{c}|_{K})$ denote the obtained signal. We always assume some underlying abelian group $G$, so we drop the notation of $G$ where it is clear. By the convolution theorem, we know that $\hat{c}_{\chi}=\sqrt{\left|G\right|}\cdot\hat{a}_{\chi}\cdot\hat{b}_{\chi}$, and so it is intuitive to think that large magnitude coefficients in $\hat{c}$ must originate from large magnitude coefficients in $\hat{a},\hat{b}$. We would have liked to say that $\left\lVert c-c^{\prime}\right\rVert_{2}\leq\left\lVert\hat{a}|_{K^{c}}\right\rVert_{2}\cdot\|\hat{b}|_{K^{c}}\|_{2}$, but unfortunately this is not correct. A correct statement would be:
Lemma 9.2.
It holds $\left\lVert c^{\prime}-c\right\rVert_{2}^{2}=\left\lVert\hat{c}|_{K}-\hat{c}\right\rVert_{2}^{2}\leq\left|G\right|\cdot\left\lVert\hat{a}|_{K^{c}}\right\rVert_{2}^{2}\|\hat{b}|_{K^{c}}\|_{2}^{2}$.
Proof.
The first equality follows from Parseval’s identity. The inequality follows from viewing $c-c^{\prime}=\mathcal{F}^{-1}(\hat{c}|_{K^{c}})$ as the convolution of two signals $a^{\prime}=\mathcal{F}^{-1}(\hat{a}|_{K^{c}})$ and $b^{\prime}=\mathcal{F}^{-1}(\hat{b}|_{K^{c}})$ (we can do this using the convolution theorem), and so by Young’s inequality (which says $\left\lVert x*y\right\rVert_{2}\leq\left\lVert x\right\rVert_{1}\cdot\left\lVert y\right\rVert_{2}$) and the $\ell^{1}$-$\ell^{2}$ inequality ($\left\lVert x\right\rVert_{1}\leq\sqrt{n}\cdot\left\lVert x\right\rVert_{2}$ for $x\in\mathbb{C}^{n}$), we obtain
$\left\lVert c-c^{\prime}\right\rVert_{2}^{2}=\left\lVert a^{\prime}*b^{\prime}\right\rVert_{2}^{2}\leq\left\lVert a^{\prime}\right\rVert_{1}^{2}\cdot\left\lVert b^{\prime}\right\rVert_{2}^{2}\leq\left|G\right|\cdot\left\lVert a^{\prime}\right\rVert_{2}^{2}\cdot\left\lVert b^{\prime}\right\rVert_{2}\overset{\text{Parseval}}{=}\left|G\right|\cdot\left\lVert\hat{a}|_{K^{c}}\right\rVert_{2}^{2}\cdot\left\lVert\hat{b}|_{K^{c}}\right\rVert_{2}^{2}.$
The $\ell^{1}$-$\ell^{2}$ is tight when the coordinates of the vector are uniform in magnitude, and loose when the coordinates are highly concentrated. So in some cases the bound might be improved upon.
The actual error in this approach is not given by the vector $c^{\prime}-c$, but rather by the projection of this vector onto the embedding set of the output matrix. Suppose $I_{C}$ denotes the indexing matrix of the output, then we care about $\left\lVert I_{C}^{\top}(c^{\prime}-c)\right\rVert_{2}^{2}$. Note that $I_{C}^{\top}$ is a contraction (because it simply drops coordinates), so the above expression is an upper bound of the error. Essentially, this shows that the error can be bounded by $\ell^{2}$-concentration of $\hat{a},\hat{b}$ on the set $K$.
Remark 9.3.
The convolution c=abc=a*b usually has large support (almost $\left|\mathrm{supp}(a)\right|\cdot\left|\mathrm{supp}(b)\right|$), but in general we cannot assume anything about the ratio $\left\lVert I_{C}^{\top}c\right\rVert_{2}/{\left\lVert c\right\rVert_{2}}$. For example, in the Vanilla TPP construction, we can take $\mathbf{A}$ to be all zeros except for $\mathbf{A}_{1,1}=M>0$ and similarly take $\mathbf{B}$ to be all zeros except for $\mathbf{B}_{2,1}=1$. Then $a$ and $b$ are all zeros except for two coordinates, whose product is not a desired term, leading to c=abc=a*b satisfying $I_{C}^{\top}c=0$ while $\left\lVert c\right\rVert_{2}=M$. Thus the ratio may be 0, or very close to it, in case there is some dominating non-desired term. In such cases, an approximation of $c$ might essentially be an approximation of the non-desired terms. This is the regime of the problem where our algorithm under performs, and is more suitable for output-sensitive algorithms, like that of [Pag13, BN20, Lin09].
The Fourier transforms for the different constructions presented here have very rigid structures:
  1. 1.
    CKSU STPP: Consider the case of $N=1$ (for $N>1$ the idea is the same, just with more notation and dimensions), i.e., the group $\mathbb{Z}_{m}^{3}$. We can write $a=a_{1}+a_{2}$ where $a_{1},a_{2}$ correspond to $\mathbf{A}_{1},\mathbf{A}_{2}$ respectively. Then $a_{1}$ is supported on $\left\{(x,y,0)\mid x,y\not=0\right\}$, which implies
    $(\hat{a}_{1})_{\xi}=C\cdot\sum_{x=1}^{m-1}\sum_{y=1}^{m-1}a_{x,y,0}\omega^{\xi_{1}x+\xi_{2}y},$
    (where $\omega=\exp(-2\pi i/m)$) showing that the spectrum of $\hat{a}_{1}$ is actually $2$-dimensional and fixed on the third axis (independent of $k_{3}$). Similarly, $a_{2}$ is supported on $\left\{(0,x,y)\right\}$ and so the spectrum of $\hat{a}_{2}$ is fixed on the first axis. The same idea is true for $b$, just with different axes.
  2. 2.
    Vanilla TPP: Recall that $a$ is supported on an arithmetic progression, with step size $n$. We have
    $\hat{a}_{\xi}=C\sum_{s=0}^{n-1}\sum_{t=0}^{n-1}a_{sn^{2}+nt}\omega^{\xi(sn^{2}+tn)},$
    (where $\omega=\exp(-2\pi i/n^{3})$) thus if $\xi_{1}\equiv\xi^{\prime}_{1}\pmod{n^{2}}$ we obtain that $\omega^{\xi_{1}(sn^{2}+tn)}=\omega^{\xi_{1}^{\prime}(sn^{2}+tn)}$ for every $s$ and $t$. Thus, $\hat{a}$ is $n^{2}$-periodic. Therefore, to estimate the top-$r$ Fourier coefficients of the convolution (assuming $r\geq n^{2}$), it suffices to estimate the top coefficients of $\hat{b}$, because we have an implicit representation of size $n^{2}$ for $\hat{a}$ which can be computed in time $O(n^{2}\log n)$. Sadly, $\hat{b}$ is not periodic. However, in the Deep Learning context, the weights matrix is fixed, and so using this construction one can treat the weights matrix (or its transpose) as $\mathbf{B}$, and compute the full spectrum of $\hat{b}$ offline. Then, assuming it is concentrated, the heavy set can be pre-computed and compared with $\hat{a}$ in real time.
9.2 Concentration In Product Distributions
This subsection is devoted to proving a concentration results for the special case in which we are given matrices whose entries are i.i.d. from a real product distribution. Throughout this subsection we will work with the CKSU STPP construction.
Notation: Let us start by fixing some notation. We are working in the group $G=\mathbb{Z}_{m}^{3}$. Define
$X_{1}=\left\{(x,y,0)\mid x,y\not=0\right\}\quad,\quad X_{2}=\left\{(0,x,y)\mid x,y\not=0\right\}\quad,\quad X_{3}=\left\{(x,0,y)\mid x,y\not=0\right\}.$
Recall that these sets form the STPP sets in the construction. In particular, given matrices $\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{B}_{1},\mathbf{B}_{2},$ we embed $\mathbf{A}_{1}$ on $X_{1}$, $\mathbf{A}_{2}$ and $\mathbf{B}_{1}$ on $X_{2}$, and $\mathbf{B}_{2}$ on $X_{3}$. Let $f_{i}=\mathbf{1}_{X_{i}}$ denote the indicator function of $X_{i}$. Our $a$ signal, is constructed by two signals $a=a_{1}+a_{2}$, such that $a_{1},a_{2}$ are encoding $\mathbf{A}_{1},\mathbf{A}_{2}$, respectively. Then $a$ will be supported on $X_{1}\cup X_{2}$, so we define $f_{a}:=\mathbf{1}_{X_{1}\cup X_{2}}=f_{1}+f_{2}$ (using the fact the set $X_{i}$ are pairwise disjoint). Similarly, $b=b_{1}+b_{2}$ will be supported on $X_{2}\cup X_{3}$, so we define $f_{b}:=f_{2}+f_{3}$. The output coordinates, which we read off to obtain our output, are $X_{1}\cup X_{3}$.
[…]
Given a real sub-Gaussian distribution $\mathcal{D}$ we denote the distribution’s parameters as:
$\mu=\mathbb{E}_{x\sim\mathcal{D}}[x],\quad\nu=\mathbb{E}_{x\sim\mathcal{D}}[x^{2}],\quad\sigma^{2}=\mathsf{Var}(\mathcal{D})=\nu-\mu^{2},$
and we let $p$ denote the variance proxy of the squared distribution, $x^{2}$ for $x\sim\mathcal{D}$. The variance proxy of a distribution $\mathcal{D}$ is the smallest real number to satisfy $\mathbb{E}_{x\sim\mathcal{D}}[e^{\lambda(x-\mathbb{E}x)}]\leq e^{p^{2}\lambda^{2}/2}$ for any $\lambda\in\mathbb{R}$. In case $\mathcal{D}$ is a bounded in $[a,b]$, $p\leq\frac{(b-a)}{2}$ (Hoeffding’s lemma). We assume all the parameters are bounded.
We will also need the following useful fact:
Fact 9.4.
Let $1<n\in\mathbb{N}$, and let $\zeta$ denote a primitive root of unity of order $n$. Then for every $k\in\mathbb{Z}$, it holds:
$\sum_{j=0}^{n-1}\zeta^{j\cdot k}=\begin{cases}0&k\not\equiv 0\pmod{n},\\ n&\text{else.}\end{cases}$
The most basic lemma we need is the spectrum structure of the indicator functions:
Lemma 9.5.
For $i\in\left\{1,2,3\right\}$ and $\xi\in\mathbb{Z}_{m}^{3}$, we have
$\hat{f}_{i}(\xi)=\frac{1}{m^{3/2}}\times\begin{cases}(m-1)^{2}&\xi_{i}=\xi_{i+1}=0,\\ -(m-1)&\xi_{i}=0\text{ xor }\xi_{i+1}=0,\\ 1&\xi_{i},\xi_{i+1}\not=0.\end{cases}$
Proof.
A direct computation shows:
$\hat{f}_{i}(\xi)$ $=\frac{1}{\sqrt{\left|\mathbb{Z}_{m}^{3}\right|}}\sum_{(x,y,z)\in\mathbb{Z}_{m}^{3}}f_{i}(x,y,z)\cdot\omega^{\xi_{1}x+\xi_{2}y+\xi_{3}z}$
$=\frac{1}{m^{3/2}}\sum_{(x,y,z)\in X_{i}}1\cdot\omega^{\xi_{1}x+\xi_{2}y+\xi_{3}z}$
$=\frac{1}{m^{3/2}}\sum_{0\not=x,y\in[m]}\omega^{\xi_{i}x+\xi_{i+1}y}$
$=\frac{1}{m^{3/2}}\left(\sum_{0\not=x\in[m]}\omega^{\xi_{i}x}\right)\left(\sum_{0\not=y\in[m]}\omega^{\xi_{i+1}y}\right)$
$=\frac{1}{m^{3/2}}\left(-1+\sum_{x\in[m]}\omega^{\xi_{i}x}\right)\left(-1+\sum_{y\in[m]}\omega^{\xi_{i+1}y}\right),$
and so the claim follows from 9.4. ∎
9.2.1 Main Theorem Statement
Now, we can state the main concentration theorem, i.e., the exact version of Theorem 1.4:
Theorem 9.6 (Concentration Theorem).
Let $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ be two real sub-Gaussian distributions, with non-zero variance and second moments. Suppose $\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{B}_{1},\mathbf{B}_{2}\in\mathbb{R}^{m-1\times m-1}$ take independent values in distributions $\mathcal{D}_{A},\mathcal{D}_{B}$. Let $r\leq m$ be a parameter, and define $K=K_{1}\cup K_{2}\cup K_{3}$ where $K_{i}=\left\{\xi\in\mathbb{Z}_{m}^{3}\mid\xi_{i}<r\right\}$. Consider the convolution of the embedded signals $a=a_{1}+a_{2},b=b_{1}+b_{2}\in\mathbb{C}[\mathbb{Z}_{m}^{3}]$ (using the CKSU STPP construction), denoted by c=abc=a*b. Let $c^{\prime}=\mathcal{F}^{-1}(\widehat{a*b}|_{K})$ denote the approximation of $c$. Then the expected normalized error of the algorithm satisfies
$\mathbb{E}\left(\frac{\|I_{C}^{\top}(c^{\prime}-a*b)\|_{2}^{2}}{\left\lVert a\right\rVert_{2}^{2}\left\lVert b\right\rVert_{2}^{2}}\right)\leq O\left(\frac{(m-r)^{3}}{(m-1)^{4}}\cdot\frac{\sigma_{A}^{2}\sigma_{B}^{2}}{\nu_{A}\nu_{B}}\right)+8(m-1)^{2}\cdot\exp(-\gamma(m-1)^{2}),$
where $\gamma=(6-4\sqrt{2})\cdot\min\left(\frac{\nu_{A}^{2}}{p_{A}^{2}},\frac{\nu_{B}^{2}}{p_{B}^{2}}\right)$.
We call the sets $K_{1},K_{2},K_{3}$, slabs of width $r$, as depicted in Figure 2.
9.2.2 Proving the Main Theorem
We start with a basic lemma, whose proof is deferred to Appendix D.
Lemma 9.7 (Low Correlation Lemma).
In the same setup of Theorem 9.6, let $\xi,\eta\in K^{c}$ and $\delta=\xi-\eta$. Then there exist functions $C_{a},C_{b}:\mathbb{Z}_{m}^{3}\to\mathbb{R}$ such that $\mathbb{E}[\hat{a}(\xi)\cdot\overline{\hat{a}(\eta)}]=C_{a}(\delta)$ and $\mathbb{E}[\hat{b}(\xi)\cdot\overline{\hat{b}(\eta)}]=C_{b}(\delta)$. Moreover, we estimate
$\left|C_{a}(\delta)\right|\leq\frac{4\mu_{A}^{2}}{m^{3}}+\sigma_{A}^{2}\cdot\begin{cases}\frac{2}{m}&\delta\in\left\{(0,0,*),(*,0,0)\right\},\\ \frac{2}{m^{2}}&\delta_{1}=0,\delta_{3}=0,\\ \frac{2}{m^{3}}&\text{otherwise,}\end{cases},\left|C_{b}(\delta)\right|\leq\frac{4\mu_{B}^{2}}{m^{3}}+\sigma_{B}^{2}\cdot\begin{cases}\frac{2}{m}&\delta\in\left\{(*,0,0),(0,*,0)\right\},\\ \frac{2}{m^{2}}&\delta_{1}=0,\delta_{2}=0,\\ \frac{2}{m^{3}}&\text{otherwise.}\end{cases}$
Notation. Let us introduce some convenient notation. Write $\mathcal{E}=\sum_{x\in X_{1}\cup X_{3}}\mathbb{E}[|e(x)|^{2}]$, where $e=c^{\prime}-a*b$ from the theorem’s statement. By the Fourier inversion formula,
$\mathcal{E}=\sum_{x\in X_{1}\cup X_{3}}\mathbb{E}\left[\left|\frac{1}{m^{3/2}}\sum_{\xi\in K^{c}}\hat{e}(\xi)\cdot\omega^{x\cdot\xi}\right|^{2}\right]=\frac{1}{m^{3}}\sum_{x\in X_{1}\cup X_{3}}\sum_{\xi,\eta\in K^{c}}\mathbb{E}[\hat{e}(\xi)\cdot\overline{\hat{e}(\eta)}]\cdot\omega^{x(\xi-\eta)}.$
Denote $\delta=\xi-\eta$. By the Convolution Theorem (formulated as $\widehat{a*b}(\xi)=\sqrt{\left|G\right|}\cdot\hat{a}(\xi)\cdot\hat{b}(\xi)$), we know $\hat{e}$ is supported on $K^{c}$, where $\hat{e}(\xi)=\widehat{a*b}(\xi)$, therefore by 9.7 we conclude that
$\mathbb{E}[\hat{e}(\xi)\cdot\overline{\hat{e}(\eta)}]=\mathbb{E}[m^{3}\cdot\hat{a}(\xi)\cdot\hat{b}(\xi)\cdot\overline{\hat{a}(\eta)\cdot\hat{b}(\eta)}]=m^{3}\cdot C_{a}(\delta)\cdot C_{b}(\delta).$
Define
$S(\delta)=\left|K^{c}\cap(K^{c}-\delta)\right|\leq\left|K^{c}\right|=(m-r)^{3}.$
Note that $S(\delta)$ is exactly the number of pairs $\xi,\eta\in K^{c}$ such that $\xi-\eta=\delta$. Thus we can change the summation order:
$\mathcal{E}=\sum_{x\in X_{1}\cup X_{3}}\sum_{\delta\in\mathbb{Z}_{m}^{3}}S(\delta)\cdot C_{a}(\delta)\cdot C_{b}(\delta)\cdot\omega^{x\cdot\delta}=\sum_{\delta\in\mathbb{Z}_{m}^{3}}S(\delta)\cdot C_{a}(\delta)\cdot C_{b}(\delta)\cdot\left(\sum_{x\in X_{1}\cup X_{3}}\omega^{x\cdot\delta}\right).$
Define $\mathcal{E}_{i}=\sum_{\delta,\left\lVert\delta\right\rVert_{0}=i}S(\delta)C_{a}(\delta)C_{b}(\delta)T(\delta)$, for $i\in\left\{0,\ldots,3\right\}$. Denote $T(\delta)=\sum_{x\in X_{1}\cup X_{3}}\omega^{x\cdot\delta}$. The next lemma gives us estimates on this quantity. The proof is deferred to Appendix D.
Lemma 9.8.
The following estimate holds:
$\left|T(\delta)\right|\leq\begin{cases}2m^{2}&\delta\in\left\{0,(0,*,0),(0,0,*)\right\},\\ 2m&\delta\in\left\{(*,0,0)\right\}\text{ or }\left\lVert\delta\right\rVert_{0}=2,\\ 2&\left\lVert\delta\right\rVert_{0}=3.\end{cases}$
Corollary 9.9.
It holds that $\mathcal{E}=\sum_{i=0}^{3}\mathcal{E}_{i}=O\left({(m-r)^{3}\cdot\sigma_{A}^{2}\cdot\sigma_{B}^{2}}\right)$.
Proof of 9.9.
Using the estimates we obtained in 9.7 and 9.8, we obtain the following table, which contains upper bounds on each quantity. For readability, we omit the $\frac{4\mu_{A}^{2}}{m^{3}}$ factor from the estimates (it is a lower order term that doesn’t change the result). Moreover, when writing (,0,0)(*,0,0) for example, we mean that $\delta_{1}\not=0$ while $\delta_{2}=\delta_{3}=0$.
$\delta$ Count $S(\delta)$ $\left|C_{a}(\delta)\right|$ $\left|C_{b}(\delta)\right|$ $\left|T(\delta)\right|$ $\textbf{Contribution to }\mathcal{E}$
0 11 $(m-r)^{3}$ $2\sigma_{A}^{2}/m$ $2\sigma_{B}^{2}/m$ $2m^{2}$ $(m-r)^{3}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}$
(,0,0)(*,0,0) (m1)(m-1) $(m-r)^{3}$ $2\sigma_{A}^{2}/m$ $2\sigma_{B}^{2}/m$ $2m$ $(m-r)^{3}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\frac{m-1}{m}$
(0,,0)(0,*,0) (m1)(m-1) $(m-r)^{3}$ $2\sigma_{A}^{2}/m^{2}$ $2\sigma_{B}^{2}/m$ $2m^{2}$ $(m-r)^{3}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\frac{m-1}{m}$
(0,0,)(0,0,*) (m1)(m-1) $(m-r)^{3}$ $2\sigma_{A}^{2}/m$ $2\sigma_{B}^{2}/m^{2}$ $2m^{2}$ $(m-r)^{3}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\frac{m-1}{m}$
$\left\lVert\delta\right\rVert_{0}=2$ $3(m-1)^{2}$ $(m-r)^{3}$ $2\sigma_{A}^{2}/m^{2}$ $2\sigma_{B}^{2}/m^{2}$ $2m$ $\frac{(m-r)^{3}}{m}\cdot 3\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\left(\frac{m-1}{m}\right)^{2}$
$\left\lVert\delta\right\rVert_{0}=3$ $(m-1)^{3}$ $(m-r)^{3}$ $2\sigma_{A}^{2}/m^{3}$ $\sigma_{B}^{2}/m^{3} 22$ $\frac{(m-r)^{3}}{m^{3}}\cdot 8\sigma_{A}^{2}\sigma_{B}^{2}\cdot\left(\frac{m-1}{m}\right)^{3}$
Using the computations done in the table above we conclude that
$\mathcal{E}=O((m-r)^{3}\cdot\sigma_{A}^{2}\cdot\sigma_{B}^{2}),$
where the constant is hiding in the $O$-notation is around $\approx 32$ (depending how big $m$ is). ∎
We will also need the following concentration bound, which is obtained by applying sub-Gaussian concentration bounds. The proof is deferred to Appendix D.
Lemma 9.10.
In the conditions and notations of Theorem 9.6, the random variable $Y=\left\lVert a\right\rVert_{2}^{2}\cdot\left\lVert b\right\rVert_{2}^{2}$ satisfies $\mathbb{E}[Y]=\nu_{A}\nu_{B}\cdot 4(m-1)^{4}$ and $\Pr(Y<\frac{1}{2}\mathbb{E}[Y])\leq 4\exp(-\gamma(m-1)^{2})$.
We are now ready to prove the theorem:
Proof of Theorem 9.6.
First we note that $\sum_{x\in X_{1}\cup X_{3}}\left|e(x)\right|^{2}=\left\lVert I_{C}^{\top}(c^{\prime}-a*b)\right\rVert_{2}^{2}$, by definition of the STPP sets. Define $X=\sum_{x\in X_{1}\cup X_{3}}\left|e(x)\right|^{2}$ and $Y=\left\lVert a\right\rVert_{2}^{2}\cdot\left\lVert b\right\rVert_{2}^{2}$. Our goal is to bound $\mathbb{E}[X/Y]$. The first step is to reduce this to a bound on $\mathbb{E}[X]/\mathbb{E}[Y]$, which is easier handle, since $\mathbb{E}[Y]=4(m-1)^{4}\cdot\nu_{A}\cdot\nu_{B}$ and $\mathbb{E}[X]=\mathcal{E}$.
By 9.10, with probability at most $4\exp(-\gamma(m-1)^{2})$ we have $Y<\frac{1}{2}\mathbb{E}[Y]$, hence
$\mathbb{E}[X/Y]\leq\mathbb{E}\left[\frac{X}{Y}\mid Y<\frac{1}{2}\mathbb{E}[Y]\right]\cdot 4\exp(-\gamma(m-1)^{2})+\frac{2\mathbb{E}[X]}{\mathbb{E}[Y]}\cdot 1.$
Since $I_{C}^{\top}$ is a contraction, we have $X\leq\left\lVert c^{\prime}-a*b\right\rVert_{2}^{2}=\left\lVert e\right\rVert_{2}^{2}=\|{\widehat{a*b}|_{K^{c}}}\|_{2}^{2}$, where we used the linearity of the Fourier transform and Parseval. Since the projection onto $K^{c}$ is also a contraction, we have $X\leq\|\widehat{a*b}\|_{2}^{2}=\left\lVert a*b\right\rVert_{2}^{2}$ which by Cauchy-Schwarz and Young’s inequality satisfies $\left\lVert a*b\right\rVert_{2}^{2}\leq 2(m-1)^{2}\cdot\left\lVert a\right\rVert_{2}^{2}\cdot\left\lVert b\right\rVert_{2}^{2}$. Therefore $X/Y\leq 2(m-1)^{2}$ always, concluding
$\mathbb{E}[X/Y]\leq 8(m-1)^{2}\cdot\exp(-\gamma(m-1)^{2})+\frac{2\mathbb{E}[X]}{\mathbb{E}[Y]},$
so we have reduced the problem to the ratio of expectations.
By 9.9 we obtain that
$\frac{2\mathbb{E}[X]}{\mathbb{E}[Y]}=\frac{\mathcal{E}}{2(m-1)^{4}\cdot\nu_{A}\cdot\nu_{B}}=\frac{O((m-r)^{3}\cdot\sigma_{A}^{2}\cdot\sigma_{B}^{2})}{2(m-1)^{4}\cdot\nu_{A}\cdot\nu_{B}}=O\left(\frac{(m-r)^{3}\cdot\sigma_{A}^{2}\cdot\sigma_{B}^{2}}{(m-1)^{4}\cdot\nu_{A}\cdot\nu_{B}}\right).$
The hidden constant is in the vicinity of $16$ (again, depending on the size of $m$, but for $m$ which is not too small, this holds). ∎
The added exponential factor is what we pay for choosing this type of analysis, but it is negligible for large enough $m$ (assuming the distribution is “nice” enough).
9.2.3 Applications of the Theorem
An immediate corollary is the application of the theorem for the Rademacher and Gaussian distributions.
Corollary 9.11.
Given random Rademacher or Gaussian matrices $\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{B}_{1},\mathbf{B}_{2}$ of size $(m-1)\times(m-1)$, and a parameter $r\leq m$, computing the set of frequencies $K$ defined in Theorem 9.6 results in matrices $\mathbf{C}_{1},\mathbf{C}_{2}$ for which
$\left\lVert\mathbf{A}_{1}\mathbf{B}_{1}-\mathbf{C}_{1}\right\rVert_{F}^{2}+\left\lVert\mathbf{A}_{2}\mathbf{B}_{2}-\mathbf{C}_{2}\right\rVert_{F}^{2}\leq O\left(\frac{(m-r)^{3}}{(m-1)^{4}}\right)\left(\left\lVert\mathbf{A}_{1}\right\rVert_{F}^{2}+\left\lVert\mathbf{A}_{2}\right\rVert_{F}^{2}\right)\left(\left\lVert\mathbf{B}_{1}\right\rVert_{F}^{2}+\left\lVert\mathbf{B}_{2}\right\rVert_{F}^{2}\right),$
in expectation.
Proof.
In both cases, we have $\mu=0$, $\nu=1$, $\sigma=1$ and $p^{2}=1$. The normalized error is thus expected to be at most
$C\cdot\left(\frac{(m-r)^{3}}{(m-1)^{4}}\right)+8\cdot(m-1)^{2}\cdot\exp\left((6-4\sqrt{2})\cdot(m-1)^{2}\right),$
where $C\approx 16$ is a global constant. ∎
We also note that $K$ has nice computation properties:
[…]

### v2

@@ 5.5 Extension to Approximate STPP
The number of trivial solutions, where $i=j=\ell$ and $s_{1}=s_{2},t_{1}=t_{2}$ and $u_{1}=u_{2}$, is exactly $M$. If $\widetilde{\rho}_{M,G}(\mathcal{I})=M$ then $\mathcal{I}$ is actually an STPP configuration for $G$.
In this setting, our algorithm outputs $\ell(\mathcal{I})$ matrices $\left\{\mathbf{C}_{i}\right\}_{i=1}^{\ell(\mathcal{I})}$, and the error of computing a convolution is $\sum_{i=1}^{\ell(\mathcal{I})}\left\lVert\mathbf{C}_{i}-\mathbf{A}_{i}\mathbf{B}_{i}\right\rVert_{F}^{2}$, which we normalize against $(\sum_{i}\left\lVert\mathbf{A}_{i}\right\rVert_{F}^{2})\cdot(\sum_{i}\left\lVert\mathbf{B}\right\rVert_{F}^{2})$. Following a similar proof to that of Theorem 5.6 (almost verbatim), we can derive the following lower bound:
Theorem 5.15.
Let $G$ be a finite group and $\mathcal{I}$ an (SIF). of size $\mu(\mathcal{I})=M$ and capacity $\kappa(\mathcal{I})$ Then the expected randomized error of implementing the algorithm in $G$ is $\Omega\left(\frac{\widetilde{\rho}_{M,G}(\mathcal{I})-M}{\kappa(\mathcal{I})}\right)$.
The size of $\tilde{\rho}$ becomes much harder to analyze when $k>1$ (note $k=1$ is the (Approximate Triple Product Quantity). case), due to the interference of different sets with each other. A very simple lower bound, is obtained by ignoring inter-set collisions. By 5.4, we obtain $\tilde{\rho}_{M,G}(\mathcal{I})\geq\left(\sum_{i=1}^{k}\left|S_{i}\right|^{2}\left|T_{i}\right|^{2}\left|U_{i}\right|^{2}\right)/\left|G\right|$. In the case $\mathcal{I}$ is just a sum of square matrix multiplications of shared size $m$, we have $\tilde{\rho}_{km^{3},G}(\mathcal{I})\geq\frac{km^{6}}{\left|G\right|}$.
In general, a more refined lower bound on $\tilde{\rho}$ is harder to obtain. We argue that the following Fourier analytic approach highlights why this problem is non-trivial. Write $\hat{X}$ to be the Fourier transform of the indicator of a set $X\subset G$. A simple application of the Fourier inversion formula leads to the identity
$\tilde{\rho}_{M,G}(\mathcal{I})=\frac{1}{\left|G\right|}\sum_{\chi\in\hat{G}}\left(\sum_{i=1}^{k}\hat{S_{i}}(\chi)\cdot\hat{T}_{i}(\chi)\right)\left(\sum_{i=1}^{k}\hat{T_{i}}(\chi)\cdot\hat{U}_{i}(\chi)\right)\left(\sum_{i=1}^{k}\hat{S_{i}}(\chi)\cdot\hat{U}_{i}(\chi)\right).$
The term corresponding to the trivial character is
$\frac{1}{\left|G\right|}\left(\sum_{i=1}^{k}\left|S_{i}\right|\cdot\left|T_{i}\right|\right)\left(\sum_{i=1}^{k}\left|T_{i}\right|\cdot\left|U_{i}\right|\right)\left(\sum_{i=1}^{k}\left|S_{i}\right|\cdot\left|U_{i}\right|\right),$
which is $\frac{kM^{2}}{\left|G\right|}=\frac{k^{3}m^{6}}{\left|G\right|}$ when $\mathcal{I}$ is $k$ square matrix multiplications of size $m$ (a factor of $k^{2}$ over the trivial lower bound). The question now becomes, how large is the rest of the sum (the non-trivial characters), and in particular, when isn’t it negative?
This question was studied in additive combinatorics, and relates to a property called Quasi-Randomness of sets (see [Bre14] and references within for background). In general, quasi-random sets have rapid mixing properties, which can be proved by showing the sum of non-trivial characters in the above decomposition is small in magnitude. Thus, as a rule of thumb, quasi-randomness implies higher $\tilde{\rho}$. Since our goal is to minimize $\tilde{\rho}$, this analysis suggests we are looking for highly structured and correlated sets, which are as far from quasi-random sets as possible (by making the sum over non-trivial characters as negative as possible). As such, finding good candidate sets is a hard and interesting problem.
6 Towards Group-Theoretic Practical AMM
Perhaps the most important and interesting question that arises from this work, is whether the STPP construction used in Theorem 1.3 and Theorem 1.2 can be made to work for real-life sized matrices (say $n=8K$, as in LLMs [Ail+25]). Although FFT is a practical algorithm, the asymptotics of the algorithms are bad. One sub-optimal feature that we point out in Theorem 1.3 is that it uses the STPP construction in a black-box manner. This inevitably requires the compressed matrices to match the asymptotics of the exact construction in Theorem 1.2. Schematically, this can be viewed as:
$\mathrm{S\&S}(\mathbf{A},\mathbf{B})=\mathrm{STPP}(\mathrm{Sketch}(\mathbf{A}),\mathrm{Sketch}(\mathbf{B})).$ (6.1)
Note that Equation 6.1 can be viewed as a special case of a low-degree approximation of the STPP polynomials (in this case, using STPP polynomials of lower degree, with linearly sketched coefficients). Another approach for degree reduction, is to instead sketch the convolution of the polynomials (e.g., via [Nak20]), i.e.,
[…]